The driver of a 2.0 × 103 kg red car traveling on the highway at 45m/s slams on his brakes to avoid striking a second yellow car in front of him, which had come to rest because of blocking ahead as shown in above Fig. After the brakes are applied, a constant friction force of 7.5 × 103 N acts on the car. Ignore air resistance.



(a) Determine the least distance should the brakes be applied to avoid a collision with the other vehicle? 



(b) If the distance between the vehicles is initially only 40.0 m, at what speed would the collision occur?



(c) Write your conclusive observations on the result obtained from this numerical. i.e. the importance of Physics in daily life.


1
Expert's answer
2021-12-07T09:53:16-0500

(a) Distance:


"d=v^2/(2a)=v^2/(2·F/m)=270\text{ m}."

(b) Speed:


"v_f=\sqrt{v^2-2xa}=\sqrt{v^2-2xF/m}=42\text{ km/h}."

(c) Physics lets us model and calculate different scenarios in life.